OpenStudy (anonymous):

put in order from least to greatest 7/2,-4,sqaure root of 7,5/4

OpenStudy (barrycarter):

Are you allowed to convert them all to decimals?

OpenStudy (anonymous):

\[\left\{-4,\frac{5}{4},\sqrt{7},\frac{7}{2}\right\} \]